# Parametric approximation as open quantum systems problem

### v2

@@ I Introduction
The parametric approximation, or constant amplitude approximation, is widely used for those modes of electromagnetic fields that have high intensity and has many applications [1, 2, 3, 4]. It can be regarded as a ”classical” approximation of the quantum field [5]. More precise conditions for the application of this approximation are discussed in Ref. [6, 7]. However, if one considers the parametric approximation as simply replacing the bosonic creation and annihilation operators by the fixed $c$-numbers, one completely neglects all quantum effects, including classical effects such as pump depletion. More importantly, there is no clear method for writing systematic corrections to the parametric approximation that take these effects into account.
One way is to move to the fully quantum regime and try to deal with the initial bosonic creation and annihilation operators without any reduction to $c$-numbers. However, using such an approach typically leads to problems that are not exactly solvable, even though they are solvable in the parametric approximation. Furthermore, if one solves such equations numerically by considering only the lowest Fock states, it becomes difficult to deal with the high-intensity regime assumed by the parametric approximation. Another way is to assume that the bosonic mode is approximately in a coherent state [5, 8, 9, 10]. In particular, this can be done in terms of the path-integral approach using the perturbative decomposition of the propagator of the system, which ultimately leads to perturbative corrections to the parametric approximation [8]. Alternatively, one can derive quantum stochastic equations in the Bogolubov-van Hove limit [11, 12] and then obtain the master equations for the system density matrix in the field for which a parametric-like approximation is assumed [9]. It is also possible to obtain the Maxwell-Bloch-type equations [10], i.e. the coupled equations both for the two-level system density matrix and the parameter of the coherent state.
@@
We consider the simplest nontrivial system, the Jaynes-Cummings model. More precisely, to account for the irreversible depletion of the external electromagnetic field, we consider the dissipative Jaynes-Cummings model [23]. If the initial states are chosen such that only a few lower Fock states of the bosonic mode are populated, then it can be solved explicitly. Moreover, the spectrum of such a model can be obtained exactly [24]. However, if one is interested in the dynamics in cases where many Fock states are populated, then the approach in Ref. [23] leads to an infinite hierarchy of coupled equations, and Ref. [24] requires finding eigenvectors of generic upper triangular matrices. Thus, even for this well-studied model, it is difficult to deal directly with the parametric regime. It is also worth noting that the methods of unitary perturbation theory, which take into account the reversible depletion of the pump [25], are not directly applicable in this case. However, in Ref. [26], it was shown that cumulant expansions for time-convolutionless master equations can be effectively applied to already open subsystems.
Note that in Refs. [27, 28, 29, 30], the equivalence conditions for the reduced dynamics of an open quantum system with two different types of environments are identified: one is a continuous bosonic environment leading to a unitary system-environment evolution, and the other is a discrete-mode bosonic environment leading to a system-mode (nonunitary) Lindbladian evolution with sums of zero-temperature dissipators. These effective discrete modes are called pseudomodes. Thus, the Jaynes-Cummings model with dissipation that we consider can be viewed not only as a result of the weak coupling derivation of the master equation for our system and some physical mode, but also as an effective master equation for our system and the pseudomode.
@@ II Jaynes-Cummings model with dissipation for approximately coherent mode
$H_{A}=$ $\omega_{A}\sigma^{+}\sigma^{-}$
@@
In this work, we consider the interaction in the rotating wave approximation, so we use a local dissipator acting only in space $B$. Our perturbative decomposition can also be applied to interaction Hamiltonians beyond the rotating wave approximation, but in that case, we will need to consider other dissipators.
@@ III.1 Free dynamics of a two level system
$\mathcal{K}_{0}\rho=-i[\omega_{A}\sigma^{+}\sigma^{-},\rho_{A}]$ (8)
@@ III.3 Dynamical Lamb shift and non-Markovian effects
Eq. (LABEL:eq:Kumulant3) can be equivalently rewritten in GKSL-like (see Ref. [51]) form
@@
The form of Eq. (LABEL:eq:Kumulant3) is significantly different from Eq. (LABEL:al:Kumulant2), so the stationary states of the second- and third-order equations are different. The complex structure of Eq. (LABEL:eq:Kumulant3) at the initial stages of evolution is also related to the effect we call the polishing of the density matrix. This term tends to zero and completely disappears at large times $\mathcal{K}_{3}(+\infty)=0$. However, this is not the case for even orders. For example, the cumulant $\mathcal{K}_{4}$ at $t\to\infty$ does not go to zero. The non-vanishing part of the fourth-order cumulant $\mathcal{K}_{4}(+\infty)$ is a correction to the unitary and dissipative parts of the second-order generator with $H^{(4)}=|g|^{4}\frac{\Delta\omega(\gamma^{2}-|\Gamma|^{2})}{|\Gamma|^{6}}\sigma^{+}\sigma^{-}$ and Lindblad operator $L^{(4)}=|g|^{2}\sqrt{\frac{\gamma(|\Gamma|^{2}-4\Delta\omega^{2})}{|\Gamma|^{6}}}\sigma^{-}$. Remark that the role of the fourth-order time-convolutionless equations in the usual open quantum systems setup is an actively discussed topic by itself [52, 53, 54, 55, 56, 57].
@@ IV Laser-induced polishing of the density matrix
During the early stages of evolution, $\mathcal{K}_{1}\neq 0$ but it tends exponentially fast to zero. We interpret the contribution from $\mathcal{K}_{1}$ to the solution of the equation as a rapid change or ”polishing” of the density matrix. In terms of Eq. (15), this leads to a renormalization of the initial conditions. Such fast renormalization is expressed by the operator $\mathcal{R}$ (see Appendix B).
@@
despite the fact that dynamics of populations and dynamics coherences of the two level system are already decoupled for such long times.
@@ V Dynamics for arbitrary initial state of bosonic mode
The hole dependence on $z$ and $\overline{z}$ is in the renormolization superoperator $\mathcal{R}$. As a result, the dynamics of a two-level system, weakly interacting with a cat state mode, will evolve according to
@@
This example also motivates to consider arbitrary quantizers [64, 65, 66, 67] instead of $|z\rangle\langle z|$ in the projector $\mathcal{P}$ along with the correspondent operator symbols instead of the Glauber–Sudarshan quasiprobablity.
@@ Acknowledgements.
The authors thank Yu. M. Belousov, Kh. Sh. Meretukov, R. Singh and A. S. Trushechkin for the fruitful discussions of the problems considered in this work. The work of A. Karasev was supported by the Foundation for the Advancement of Theoretical Physics and Mathematics “BASIS” under the grant № 23-2-9-19-1.
@@ A.3 Parametric approximation
Starting from this stage we will use the explicit form of $\mathcal{X}^{(1)}$, even though we could do without it in the first order. For consistency, we will calculate the cumulants after integration of all expressions for the moments. We number all $6$ elements of $\mathcal{X}^{(1)}$ in the same order as they are in the sum Eq. (A.2). It is convenient for us to work with the elements of $\mathcal{X}^{(1)}$ separately.
@@
The index $1$ denotes the part of an expression associated with $\mathcal{X}^{(1)}_{1}$. Summing up similar expressions for all indices we obtain Eq. (10).
@@ A.4 Dissipation
which is a part of Eq. (LABEL:al:Kumulant2).
@@ Appendix B Asymptotic matching
Eq. (57) satisfies the conditions of Theorem 3 from Ref. [26]. Using this Theorem, we can find the correct solution Eq. (17) of Eq. (60) (here we apply the limit of large times $t\to t/\lambda^{2}$, $\lambda\to 0$, $t/\lambda^{2}=const$).
@@ Appendix C Quasiprobability distribution
According to Ref. [69] the corresponding quasiprobability function of the Schrödinger cat state is as follows
@@
So the transformation Eq. (20) is obtained by calculating integrals $\int P(z)zd^{2}z$ and $\int P(z)\overline{z}d^{2}z$.

### v3

@@ I Introduction
The parametric approximation, or constant amplitude approximation, is widely used for those modes of electromagnetic fields that have high intensity and has many applications [1, 2, 3, 4]. It can be regarded as a ”classical” approximation of the quantum field [5]. More precise conditions for the application of this approximation are discussed in [6, 7]. However, if one considers the parametric approximation as simply replacing the bosonic creation and annihilation operators by the fixed $c$-numbers, one completely neglects all quantum effects, including classical effects such as pump depletion. More importantly, there is no clear method for writing systematic corrections to the parametric approximation that take these effects into account.
One way is to move to a fully quantum regime and try to deal with the initial bosonic creation and annihilation operators without any reduction to $c$-numbers. However, using such an approach typically leads to problems that are not exactly solvable, even though they are solvable in the parametric approximation. Furthermore, if one solves such equations numerically by considering only the lowest Fock states, it becomes difficult to handle the high-intensity regime assumed by the parametric approximation. Another way is to assume that the bosonic mode is approximately in a coherent state [5, 8, 9, 10]. In particular, this can be done in terms of the path-integral approach using the perturbative decomposition of the propagator of the system, which ultimately leads to perturbative corrections to the parametric approximation [8]. Alternatively, one can derive quantum stochastic equations in the Bogolubov-van Hove limit [11, 12] and then obtain the master equations for the system density matrix in the field for which a parametric-like approximation is assumed [9]. It is also possible to obtain the Maxwell-Bloch-type equations [10], i.e. the coupled equations both for the two-level system density matrix and the parameter of the coherent state.
@@
We consider the simplest nontrivial system, the Jaynes-Cummings model. More precisely, to account for the irreversible depletion of the external electromagnetic field, we consider the dissipative Jaynes-Cummings model [23]. If the initial states are chosen such that only a few lower Fock states of the bosonic mode are populated, then it can be solved explicitly. Moreover, the spectrum of such a model can be obtained exactly [24]. However, if one is interested in the dynamics in cases where many Fock states are populated, then the approach in [23] leads to an infinite hierarchy of coupled equations, and  [24] requires finding eigenvectors of generic upper triangular matrices. Thus, even for this well-studied model, it is difficult to deal directly with the parametric regime. It is also worth noting that the methods of unitary perturbation theory, which take into account the reversible depletion of the pump [25], are not directly applicable in this case. However, in [26], it was shown that cumulant expansions for time-convolutionless master equations can be effectively applied to already open subsystems.
Note that in [27, 28, 29, 30], the equivalence conditions for the reduced dynamics of an open quantum system with two different types of environments are identified: one is a continuous bosonic environment leading to a unitary system-environment evolution, and the other is a discrete-mode bosonic environment leading to a system-mode (nonunitary) Lindbladian evolution with sums of zero-temperature dissipators. These effective discrete modes are called pseudomodes. Thus, the Jaynes-Cummings model with dissipation that we consider can be viewed not only as a result of the weak coupling derivation of the master equation for our system and some physical mode, but also as an effective master equation for our system and the pseudomode.
@@ II Jaynes-Cummings model with dissipation for approximately coherent mode
$H_{A}=$ $\omega_{A}\sigma^{+}\sigma^{-}{,}$
@@
In this work, we consider the interaction in the rotating wave approximation, so we use a local dissipator acting only on the subsystem $B$. Our perturbative decomposition can also be applied to interaction Hamiltonians beyond the rotating wave approximation, but in that case, we will need to consider other dissipators.
@@ III.1 Free dynamics of a two level system
$\mathcal{K}_{0}\rho=-i[\omega_{A}\sigma^{+}\sigma^{-},\rho_{A}]{.}$ (8)
@@ III.3 Dynamical Lamb shift and non-Markovian effects
The most direct effect of field depletion on the experimentally observable quantities is the change in the Mollow triplet. A ”naive” approach would be to substitute the instantaneous depleted field in the usual formula for the Mollow spectrum for the undepleted field. However, such an approach completely ignores all quantum properties of the field. We can compare the results obtained with our quantum treatment (see Appendix E for details). Then we obtain a more squeezed but also more contrasted picture (in terms of the ratio of side peaks to intermediate minima) of the Mollow triplet, which is shown in Fig. 1.
Figure 1: The Mollow triplet based on classical and quantum treatment is shown for the specific choice of parameters $\lambda=0.4$, $g=\overline{g}=\gamma$, $z=\overline{z}=7\gamma$, $\Delta\omega=0$, $\gamma t=2$. The dashed blue line corresponds to the ”naive” substitution of the depleted field into the formula for the undepleted case, ignoring all quantum effects. The solid red line corresponds to the quantum master equations obtained in this work. The ”naive” treatment also has a delta function contribution in the center, which is omitted from the plot.
@@
(LABEL:eq:Kumulant3) can be equivalently rewritten in GKSL-like (see [51]) form
@@
Remark that the third-order contribution to the dissipator couples the populations and coherences, in contrast to (LABEL:al:Kumulant2), which only leads to a decoupled classical Markovian dynamics for populations. In this sense, the second-order contribution captures only classical dissipation effects, while the higher-order contributions essentially account for quantum dissipation effects. The higher-order terms also affect the instantaneous steady state of the dissipator, which is important for applications of master equations in quantum thermodynamics [52].
The form of (LABEL:eq:Kumulant3) is significantly different from (LABEL:al:Kumulant2), so the stationary states of the second- and third-order equations are different. The complex structure of (LABEL:eq:Kumulant3) at the initial stages of evolution is also related to the effect we call the polishing of the density matrix. This term tends to zero and completely disappears at large times $\mathcal{K}_{3}(+\infty)=0$. However, this is not the case for even orders. For example, the cumulant $\mathcal{K}_{4}$ at $t\to\infty$ does not go to zero. The non-vanishing part of the fourth-order cumulant $\mathcal{K}_{4}(+\infty)$ is a correction to the unitary and dissipative parts of the second-order generator with $H^{(4)}=|g|^{4}\frac{\Delta\omega(\gamma^{2}-|\Gamma|^{2})}{|\Gamma|^{6}}\sigma^{+}\sigma^{-}$ and Lindblad operator $L^{(4)}=|g|^{2}\sqrt{\frac{\gamma||\Gamma|^{2}-4\Delta\omega^{2}|}{|\Gamma|^{6}}}\sigma^{-}$, which define the Lindblad contribution for $|\Gamma|^{2}-4\Delta\omega^{2}\geq 0$ and the Lindblad-like contribution with a negative sing for $|\Gamma|^{2}-4\Delta\omega^{2}<0$. Thus, this term lowers the long-time population and coherence decay rates in the significantly off-resonance case. Remark that the role of the fourth-order time-convolutionless equations in the usual open quantum systems setup is an actively discussed topic by itself [53, 54, 55, 56, 57, 58].
@@ IV Laser-induced polishing of the density matrix
Qualitatively, we have the following picture. The interaction with the depleting field goes through two stages. In the first stage, the coherent interaction dominates. In the second stage, the field is almost completely depleted and the non-coherent interaction is dominant, but the effect of the first stage still contributes to the dynamics in the second stage by preparing the density matrix for the start of the second stage. We call this preparation ”laser-induced polishing”. In terms of the master equation at the first stage $\mathcal{K}_{1}\neq 0$ but it tends exponentially fast to zero. And in the second state the density matrix satisfies (15), but the ”laser-induced polishing” is manifested in the renormalization of the initial conditions. Such renormalization is expressed by the operator $\mathcal{R}$ (see Appendix B).
@@
despite the fact that dynamics of populations and coherences of the two level system are already decoupled for such long times.
Figure 2: All three graphs show the average $\langle\sigma_{+}\sigma_{-}(t)\rangle$ as a function of $\gamma t$, i.e. of rescaled time in the units of the bosonic mode decay time, for a specific choice of parameters $\lambda=0.055$, $g=\overline{g}=20\gamma$, $z=\overline{z}=342\gamma$, $\Delta\omega=1.3\gamma$. $\rho(0)$ is the eigenstate of $\sigma^{x}$ corresponding to the eigenvalue $+1$. The solid blue line is obtained by the time-dependent generator. The dashed red line is obtained by the long-time equation (15) with renormalized initial condition (16). And dot-dashed green line is obtained by the long-time equation (15) with the initial condition taken without renormalization. The initial oscillatory behavior of the solid blue line corresponds to the polishing of the density matrix during the first stage. The decaying behavior of the solid blue line corresponds, where it is close to dashed red line, corresponds to the second stage, at the beginning of which the density matrix is already polished. However, without taking into account the polishing via renormalization of the initial density matrix, the long-time equations cannot correctly reproduce dynamics at the second stage, which is reflected in the fact that the dot-dashed green line is far from solid blue line during most of the dynamics.
We illustrate the physical picture discussed in this section in Fig. 2, where $\langle\sigma_{-}\sigma_{+}\rangle(t)$ as a function of time is presented, since it is proportional to the intensity of the radiated field [59, Eq. (15.6.3)] which is experimentally measurable. Namely, we present the prediction of time-dependent second order master equations together with the long-time master equation (15) with and without renormalization. A similar picture with a clear separation of two stages will occur under the condition that the decay rate in (17) is smaller than the depletion rate in (10), but the field is strong enough to have a non-negligible contribution in (16), i.e. if
$|g|\lesssim|\Gamma|,\qquad\lambda|g||z|\sim|\Gamma|.$ (18)
Remark, that we choose the parameters such that coherent oscillations are clearly visible in Fig. 2, but for higher intensity of the initial input field they can be much faster, so the form from the experimental point of view they will look like stochastic behavior around some averaged dynamics, while the second stage has more regular behavior. In this case, the laser-induced polishing will look like preparation of the density matrix during some stochastic behavior for the more regular dynamics.
@@ V Dynamics for arbitrary initial state of bosonic mode
The whole dependence on $z$ and $\overline{z}$ is in the renormalization superoperator $\mathcal{R}$. As a result, the dynamics of a two-level system, weakly interacting with a cat state mode, will evolve according to
@@
This example also motivates to consider arbitrary quantizers [66, 67, 68, 69] instead of $|z\rangle\langle z|$ in the projector $\mathcal{P}$ along with the correspondent operator symbols instead of the Glauber–Sudarshan quasiprobablity. This generalization to other quantizers can also be used to extend our approach to other types of environmental degrees of freedom instead of the bosonic mode, e.g. to tensor products of qubits or oscillating spin-1/2 particles. It seems to be particularly effective for those states for which their probability representation [70] is explicitly known together with the corresponding quantizers, e.g. for cat states of two or three qubits, or of an oscillating spin-1/2 particle [71].
@@ Acknowledgements.
The authors thank Yu. M. Belousov, Kh. Sh. Meretukov, R. Singh and A. S. Trushechkin for the fruitful discussions of the problems considered in this work. The authors thank the reviewers for their comments, which led to significant improvement of the paper. The work of A. Karasev was supported by the Foundation for the Advancement of Theoretical Physics and Mathematics “BASIS” under the grant № 24-2-1-48-1.
@@ A.3 Parametric approximation
Starting from this stage we will use the explicit form of $\mathcal{X}^{(1)}$, even though we could do without it in the first order. For consistency, we will calculate the cumulants after integration of all expressions for the moments. We number all $6$ elements of $\mathcal{X}^{(1)}$ in the same order as they are in the sum (A.2). It is convenient for us to work with the elements of $\mathcal{X}^{(1)}$ separately.
@@
The index $1$ denotes the part of an expression associated with $\mathcal{X}^{(1)}_{1}$. Summing up similar expressions for all indices we obtain (10).
@@ A.4 Dissipation
which is a part of (LABEL:al:Kumulant2).
@@ Appendix B Asymptotic matching
(58) satisfies the conditions of Theorem 3 from [26]. Using this Theorem, we can find the correct solution (17) of (61) (here we apply the limit of large times $t\to t/\lambda^{2}$, $\lambda\to 0$, $t/\lambda^{2}=const$).
@@ Appendix C Quasiprobability distribution
According to [73] the corresponding quasiprobability function of the Schrödinger cat state is as follows
@@
So the transformation (21) is obtained by calculating integrals $\int P(z)zd^{2}z$ and $\int P(z)\overline{z}d^{2}z$.
@@ Appendix D Vectorization
Appendix E Resonance fluorescence spectrum for depleting field
The resonance fluorescence spectrum is proportional to the function [76, Eqs. (10.5.1)-(10.5.2)]
$S(\omega)=\operatorname{Re}\int_{0}^{\infty}\langle\sigma_{+}(t)\sigma_{-}(t+\tau)\rangle e^{i\omega\tau}.$ (90)
Assuming that the total dynamics of the atom and the mode of the field is Markovian it can be defined by the quantum regression formula [13, Subsec. 3.2.4]
$\langle\sigma_{+}(t)\sigma_{-}(t+\tau)\rangle=\operatorname{Tr}_{S+B}\mathcal{U}_{\tau}(\sigma_{+}\rho_{\rm st}(t))\sigma_{-},$ (91)
where $\tau>0$, $t>0$ and $\mathcal{U}_{\tau}$ is the total dynamical map of the atom and the bosonic mode. For the calculation of the Mollow spectrum in the undepleted field, $\rho_{\rm st}(t)$ is usually assumed to be in the stationary state and factorized so such that the quantum regression formula can be applied to the reduced dynamics of the atom [13, Subsec. 3.4.5]. Thus, we use a natural generalization to the case of depleting filed, assuming that in the approximately stationary state is also factorized
$\rho_{\rm st}(t)\approx(\rho_{A})_{\rm st}(t)\otimes|z(t)\rangle\langle z(t)|,$ (92)
where $(\rho_{A})_{\rm st}(t)$ is an instantaneous stationary state of the generator we are considering, and the field is approximately in the coherent state with $z(t)=ze^{(-i\omega_{c}-\frac{\gamma}{2})t}$, consistent with our assumptions for the atomic master equation derivation
Under assumption (92) correlation function (91) takes the form
$\langle\sigma_{+}(t)\sigma_{-}(t+\tau)\rangle=\operatorname{Tr}_{S}\Phi_{\tau}|_{z(t)}(\sigma_{+}(\rho_{A})_{\rm st}(t))\sigma_{-},$
where $\Phi_{\tau}|_{z(t))}$ is the dynamical map with generator $\mathcal{K}_{\tau}$ depending on time $\tau$ with $z(t)$ instead of $z$ as a parameter, so we can apply our perturbative formulae for $\mathcal{K}_{\tau}$.